# ButterflyQuant: Ultra-low-bit LLM Quantization through Learnable Orthogonal Butterfly Transforms

### v2

@@ Abstract
Large language models require massive memory footprints, severely limiting deployment on consumer hardware. Quantization reduces memory through lower numerical precision, but extreme 2-bit quantization suffers from catastrophic performance loss due to outliers in activations. Rotation-based methods such as QuIP and QuaRot apply orthogonal transforms to eliminate outliers before quantization, using computational invariance: $\mathbf{y}=\mathbf{Wx}=(\mathbf{WQ}^{T})(\mathbf{Qx})$ for orthogonal $\mathbf{Q}$. However, these methods use fixed transforms–Hadamard matrices achieving optimal worst-case coherence $\mu=1/\sqrt{n}$–that cannot adapt to specific weight distributions. We identify that different transformer layers exhibit distinct outlier patterns, motivating layer-adaptive rotations rather than one-size-fits-all approaches. In this work, we propose ButterflyQuant, which replaces Hadamard rotations with learnable butterfly transforms parameterized by continuous Givens rotation angles. Unlike Hadamard’s discrete $\{+1,-1\}$ entries that are non-differentiable and thus prohibit gradient-based learning, butterfly transforms’ continuous parameterization enables smooth optimization while guaranteeing orthogonality by construction. This orthogonal constraint ensures theoretical guarantees in outlier suppression while achieving $O(n\log n)$ computational complexity with only $\frac{n\log n}{2}$ learnable parameters. We further introduce a uniformity regularization on post-transformation activations to promote smoother distributions amenable to quantization. Learning requires only 128 calibration samples and converges in minutes on a single GPU–a negligible one-time cost. For LLaMA-2-7B with 2-bit quantization, ButterflyQuant achieves 15.4 perplexity versus 37.3 for QuIP. Codes are available.
@@ 1 Introduction
To mitigate the outlier problem, rotation-based quantization methods have emerged as a robust solution [Ashkboos et al., 2024, Chee et al., 2023]. These methods apply an orthogonal transformation $\mathbf{Q}$ before quantization, leveraging computational invariance: $\mathbf{y}=\mathbf{Wx}=(\mathbf{WQ}^{T})(\mathbf{Qx})$. The rotation redistributes activations across channels, effectively smoothing out outlier features without altering the layer’s output. Prominent methods like QuaRot [Ashkboos et al., 2024] use fixed Hadamard transforms, which achieve optimal worst-case coherence, while QuIP [Chee et al., 2023] employs random orthogonal matrices. Despite their success, these approaches share a critical limitation: they apply a fixed, data-agnostic transform with discrete $\{+1,-1\}$ entries that cannot be optimized via gradients to all layers.
@@
This one-size-fits-all strategy is fundamentally misaligned with the nature of LLMs, which exhibit significant heterogeneity across layers [Sun et al., 2024]. As illustrated in Figure 1(a), different transformer layers present unique quantization challenges: attention layers develop outliers in positive tails [Bondarenko et al., 2023], early MLP layers show them in negative regions [Wei et al., 2022], and late MLP layers have them near distribution boundaries [Sun et al., 2024]. These distinct patterns arise from varied architectural roles. Attention’s softmax operation naturally produces positive-skewed distributions [Bondarenko et al., 2023], early MLP gating functions (e.g., SwiGLU [Shazeer, 2020]) create asymmetric negative activations, while deeper layers accumulate numerical artifacts at distribution extremes [Sun et al., 2024, Dettmers et al., 2022]. This heterogeneity means that a single fixed rotation cannot be optimal for all layers. This layer-specific structure reveals a missed opportunity for optimization, motivating a shift from fixed to adaptive rotations.
@@
To address this need for an adaptive yet efficient orthogonal transform, we propose ButterflyQuant. Our method replaces fixed Hadamard rotations with learnable butterfly transforms–structured orthogonal matrices factorized into $O(n\log n)$ Givens rotations [Dao et al., 2019]. As shown in Figure 1(c), butterfly transforms maintain the efficient computational structure of Hadamard matrices but learn layer-specific rotation angles via gradient descent. Crucially, butterfly transforms parameterize rotations through continuous angles $\theta\in\mathbb{R}$, enabling smooth gradient flow and stable optimization—in contrast to Hadamard’s discrete $\{+1,-1\}$ entries that prohibit gradient-based learning. This allows them to develop adaptive coherence patterns that are tailored to each layer’s unique outlier distribution. Unlike other learnable methods (e.g., SpinQuant [Liu et al., 2024a]) that optimize over the full Stiefel manifold with high computational cost, our sparse parameterization guarantees orthogonality by construction, enabling stable and efficient optimization. For non-power-of-2 dimensions common in LLMs (e.g., 5120), we develop composite transforms using Kronecker products. The learning process is remarkably lightweight, converging in minutes on a single GPU with a small calibration set.
@@ 2.1 Post-Training Quantization for LLMs
Recent surveys [Zhao et al., 2025] categorize LLM PTQ methods into four main approaches: compensation, rotation, salience, and optimization-based techniques.
Compensation and Salience Methods. GPTQ [Frantar et al., 2023] uses second-order Hessian information for reconstruction error minimization. AWQ [Lin et al., 2024a] preserves salient weights while quantizing others based on activation magnitudes. SmoothQuant [Xiao et al., 2023] migrates difficulty from activations to weights via channel-wise scaling. OmniQuant [Shao et al., 2023] and AffineQuant [Guan et al., 2024a] apply learnable transformations for outlier suppression. Mixed-precision methods like LLM.int8() [Dettmers et al., 2022] and SpQR [Dettmers et al., 2023] maintain outlier channels in higher precision but increase complexity.
Optimization Methods. Recent approaches include VPTQ [Liu et al., 2024b] using vector quantization with optimized codebooks, APTQ [Guan et al., 2024b] with adaptive precision allocation, and AQLM [Egiazarian et al., 2024] employing additive quantization for 2-bit compression. While achieving strong results, these methods require extensive optimization and lack theoretical guarantees of rotation-based approaches.
@@ 2.2 Rotation-Based Quantization
Fixed Methods. QuaRot [Ashkboos et al., 2024] applies Hadamard transforms to redistribute outliers. QuIP [Chee et al., 2023] and QuIP# [Tseng et al., 2024] use random orthogonal matrices based on the incoherence principle, showing that maximizing incoherence minimizes worst-case quantization error. However, these predetermined rotations cannot adapt to specific models.
Learned Methods. SpinQuant [Liu et al., 2024a] optimizes full $n\times n$ rotation matrices on the Stiefel manifold, requiring $O(n^{2})$ parameters (16.7M for 4096-dim layers) and $O(n^{3})$ operations. ROSAQ [Kim et al., 2024] and KurTail [Akhondzadeh et al., 2025] learn saliency-aware and kurtosis-guided rotations respectively, but lack theoretical guarantees and still require $O(n^{2})$ space. DuQuant Lin et al. [2024b] employs a greedy construction method that uses identified outlier dimensions as prior knowledge to build rotation matrices, followed by zigzag permutation to balance outlier distribution across blocks.
Our Position. ButterflyQuant bridges fixed and learned approaches through structured parameterization. Unlike QuIP#’s fixed rotations or SpinQuant’s expensive Stiefel optimization, our Givens parameterization guarantees orthogonality by construction while enabling efficient data-driven learning, combining theoretical guarantees with adaptability.
@@ 2.3 Structured Transforms and Orthogonal Parameterization
Despite extensive use in deep learning, structured transforms remain unexplored for quantization, which requires: (1) strict orthogonality for computational invariance $\mathbf{y}=\mathbf{Wx}=(\mathbf{WQ}^{T})(\mathbf{Qx})$, (2) adaptability to layer-specific outliers [Sun et al., 2024], and (3) efficient inference. While FlatQuant [Liu et al., 2024c] uses Kronecker decomposition, it sacrifices orthogonality. ButterflyQuant uniquely combines continuous parameterization ($\theta\in\mathbb{R}$ enabling gradients vs. Hadamard’s discrete $\{+1,-1\}$) and $O(n\log n)$ complexity, balancing expressiveness with efficiency. ButterflyQuant is the pioneer in leveraging these properties for LLM quantization.
@@ 3.1 Preliminaries: Rotation-Based Quantization and Incoherence
Given a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ and activation $\mathbf{x}\in\mathbb{R}^{n}$, standard quantization directly quantizes:
@@ 3.1.1 Theoretical Foundation: The Incoherence Principle
The key insight from QuIP [Chee et al., 2023] is that quantization error is minimized when the rotation basis is maximally incoherent with the standard basis. For an orthogonal matrix $\mathbf{Q}\in\mathbb{R}^{n\times n}$, the mutual coherence is defined as:
@@ 3.1.2 Fixed Hadamard Transforms
QuaRot [Ashkboos et al., 2024] applies Hadamard transforms as a computationally efficient approximation to random rotations. The Hadamard matrix $\mathbf{H}_{n}$ of dimension $n\times n$ is recursively defined as:
@@ 3.1.3 From Fixed to Learnable Continuous Rotations
Although Hadamard transforms achieve optimal worst-case incoherence, they suffer from two critical limitations: (1) Their discrete $\{+1,-1\}$ entries prohibit gradient-based optimization, and (2) neural networks exhibit layer-specific structured patterns [Sun et al., 2024]. As shown in Figure 1(a), attention layers, early MLPs, and late MLPs each have distinct outlier distributions that require tailored rotations. Fixed transforms cannot adapt to these heterogeneous patterns, treating all layers identically despite their vastly different quantization challenges.
@@ 3.2.1 Structure and Parameterization
This parameterization ensures orthogonality by construction while enabling gradient-based optimization. See Appendix B for detailed matrix structures and visualizations.
@@ 3.2.2 Relationship to Hadamard Transforms
Theorem 1.
The Hadamard matrix $\mathbf{H}_{n}$ for $n=2^{k}$ can be exactly represented as a butterfly transform with specific parameter choices [Dao et al., 2019].
@@
$\mathbf{H}_{n}=\begin{bmatrix}\mathbf{H}_{n/2}&\mathbf{H}_{n/2}\\ \mathbf{H}_{n/2}&-\mathbf{H}_{n/2}\end{bmatrix}=\begin{bmatrix}\mathbf{I}_{n/2}&\mathbf{I}_{n/2}\\ \mathbf{I}_{n/2}&-\mathbf{I}_{n/2}\end{bmatrix}\begin{bmatrix}\mathbf{H}_{n/2}&\mathbf{0}\\ \mathbf{0}&\mathbf{H}_{n/2}\end{bmatrix}$ (10)
@@
which corresponds to a Givens rotation with $\theta=\pi/4$ (after appropriate scaling and permutation). The complete recursive factorization yields $\log_{2}n$ sparse matrices, each implementable as a butterfly layer with specific angle choices, establishing that Hadamard transforms are a special case of butterfly parameterizations. ∎
@@ 3.2.3 Theoretical Coherence Analysis
For orthogonal transforms, coherence $\mu(\mathbf{Q})=\max_{i\neq j}|Q_{ij}|$ measures how evenly information is distributed. From compressed sensing [Candès et al., 2006, Donoho, 2006], lower coherence reduces the sampling requirements for successful recovery:
@@
Figure 2 compares coherence across LLaMA-2-7B layers ($n=4096$). Hadamard transforms achieve the Welch bound $\mu(\mathbf{H}_{n})=1/\sqrt{n}=1.56\times 10^{-2}$ uniformly. Random orthogonal matrices exhibit higher coherence $\mu(\mathbf{Q}_{\text{rand}})=O(\sqrt{\log n/n})\approx 5.4\times 10^{-2}$ [Vershynin, 2018]. Learned butterfly transforms demonstrate adaptive coherence varying from $1.8$ to $3.2\times 10^{-2}$ across layers, matching the heterogeneous outlier patterns: early attention layers maintain near-Welch-bound coherence for uniform decorrelation, while deeper MLP layers relax this constraint for their specific activation patterns.
Empirical Coherence Adaptation.
The learned butterfly transforms demonstrate adaptive coherence that varies from $1.8$ to $3.2\times 10^{-2}$ across layers, matching the heterogeneous outlier patterns identified in our analysis. This adaptation—tailoring coherence to layer-specific requirements rather than enforcing uniform worst-case optimality—enables superior quantization performance.
The Givens parameterization ensures smooth optimization with stable convergence, typically achieving 80% of the improvement within 100 iterations, without the oscillations common in full matrix optimization. This efficiency, combined with the theoretical guarantees of orthogonality, makes butterfly transforms an ideal bridge between fixed and fully learnable rotations.

### v3

@@ Abstract
Large language models require massive memory footprints, severely limiting deployment on consumer hardware. Quantization reduces memory through lower numerical precision, but extreme 2-bit quantization suffers from catastrophic performance loss due to outliers in activations. Rotation-based methods such as QuIP and QuaRot apply orthogonal transforms to eliminate outliers before quantization, using computational invariance: $\mathbf{y}=\mathbf{Wx}=(\mathbf{WQ}^{T})(\mathbf{Qx})$ for orthogonal $\mathbf{Q}$. However, these methods use fixed transforms–Hadamard matrices achieving optimal worst-case coherence $\mu=1/\sqrt{n}$–that cannot adapt to specific weight distributions. We identify that different transformer layers exhibit distinct outlier patterns, motivating layer-adaptive rotations rather than one-size-fits-all approaches. In this work, we propose ButterflyQuant, which replaces Hadamard rotations with learnable butterfly transforms parameterized by continuous Givens rotation angles. Unlike Hadamard’s discrete $\{+1,-1\}$ entries that are non-differentiable and thus prohibit gradient-based learning, butterfly transforms’ continuous parameterization enables smooth optimization while guaranteeing orthogonality by construction. This orthogonal constraint ensures theoretical guarantees in outlier suppression while achieving $O(n\log n)$ computational complexity with only $\frac{n\log n}{2}$ learnable parameters. We further introduce a uniformity regularization on post-transformation activations to promote smoother distributions amenable to quantization. Learning requires only 128 calibration samples and converges in minutes on a single GPU.
Machine Learning, ICML
@@ 1 Introduction
To mitigate the outlier problem, rotation-based quantization methods have emerged as a robust solution (Ashkboos et al., 2024; Chee et al., 2023). These methods apply an orthogonal transformation $\mathbf{Q}$ before quantization, leveraging computational invariance: $\mathbf{y}=\mathbf{Wx}=(\mathbf{WQ}^{T})(\mathbf{Qx})$. The rotation redistributes activations across channels, effectively smoothing out outlier features without altering the layer’s output. Prominent methods like QuaRot (Ashkboos et al., 2024) use fixed Hadamard transforms, which achieve optimal worst-case coherence, while QuIP (Chee et al., 2023) employs random orthogonal matrices. Despite their success, these approaches share a critical limitation: they apply a fixed, data-agnostic transform with discrete $\{+1,-1\}$ entries that cannot be optimized via gradients to all layers.
@@
This one-size-fits-all strategy is fundamentally misaligned with the nature of LLMs, which exhibit significant heterogeneity across layers (Sun et al., 2024a). As illustrated in Figure 1(a), different transformer layers present unique quantization challenges: attention layers develop outliers in positive tails (Bondarenko et al., 2023), early MLP layers show them in negative regions (Wei et al., 2022), and late MLP layers have them near distribution boundaries (Sun et al., 2024a). These distinct patterns arise from varied architectural roles. Attention’s softmax operation naturally produces positive-skewed distributions (Bondarenko et al., 2023), early MLP gating functions (e.g., SwiGLU (Shazeer, 2020)) create asymmetric negative activations, while deeper layers accumulate numerical artifacts at distribution extremes (Sun et al., 2024a; Dettmers et al., 2022). This heterogeneity means that a single fixed rotation cannot be optimal for all layers. This layer-specific structure reveals a missed opportunity for optimization, motivating a shift from fixed to adaptive rotations.
@@
To address this need for an adaptive yet efficient orthogonal transform, we propose ButterflyQuant. Our method replaces fixed Hadamard rotations with learnable butterfly transforms–structured orthogonal matrices factorized into $O(n\log n)$ Givens rotations (Dao et al., 2019). As shown in Figure 1(c), butterfly transforms maintain the efficient computational structure of Hadamard matrices but learn layer-specific rotation angles via gradient descent. Crucially, butterfly transforms parameterize rotations through continuous angles $\theta\in\mathbb{R}$, enabling smooth gradient flow and stable optimization—in contrast to Hadamard’s discrete $\{+1,-1\}$ entries that prohibit gradient-based learning. This allows them to develop adaptive coherence patterns that are tailored to each layer’s unique outlier distribution. Unlike other learnable methods (e.g., SpinQuant (Liu et al., 2024c)) that optimize over the full Stiefel manifold with high computational cost, our sparse parameterization guarantees orthogonality by construction, enabling stable and efficient optimization. For non-power-of-2 dimensions common in LLMs (e.g., 5120), we develop composite transforms using Kronecker products. The learning process is remarkably lightweight, converging in minutes on a single GPU with a small calibration set.
Our contributions are: (1) We introduce learnable butterfly transforms for LLM quantization, maintaining orthogonality guarantees while adapting to model- and layer-specific weight distributions. (2) We prove that butterfly transforms can represent Hadamard matrices exactly, making them strictly more expressive than fixed schemes. (3) We achieve state-of-the-art 2-bit quantization results among rotation-based methods. (4) We show that the learned butterfly transform preserves the same $O(n\log n)$ butterfly computational graph as the fast Hadamard transform, enabling straightforward adaptation of existing optimized Hadamard CUDA kernels with expected minimal inference overhead.
@@ 2.1 Post-Training Quantization for LLMs
Recent surveys (Zhao et al., 2025) categorize LLM PTQ methods into four main approaches: compensation, rotation, salience, and optimization-based techniques.
Compensation and Salience Methods. GPTQ (Frantar et al., 2023) uses second-order Hessian information for reconstruction error minimization. AWQ (Lin et al., 2024b) preserves salient weights while quantizing others based on activation magnitudes. SmoothQuant (Xiao et al., 2023) migrates difficulty from activations to weights via channel-wise scaling. OmniQuant (Shao et al., 2023) and AffineQuant (Ma et al., 2024) apply learnable transformations for outlier suppression. Mixed-precision methods like LLM.int8() (Dettmers et al., 2022) and SpQR (Dettmers et al., 2023) maintain outlier channels in higher precision but increase complexity.
Optimization Methods. Recent approaches include VPTQ (Liu et al., 2024b) using vector quantization with optimized codebooks, APTQ (Guan et al., 2024) with adaptive precision allocation, and AQLM (Egiazarian et al., 2024) employing additive quantization for 2-bit compression. While achieving strong results, these methods require extensive optimization and lack theoretical guarantees of rotation-based approaches.
@@ 2.2 Rotation-Based Quantization
Fixed Methods. QuaRot (Ashkboos et al., 2024) applies Hadamard transforms to redistribute outliers. QuIP (Chee et al., 2023) and QuIP# (Tseng et al., 2024) use random orthogonal matrices based on the incoherence principle, showing that maximizing incoherence minimizes worst-case quantization error. However, these predetermined rotations cannot adapt to specific models.
Learned Methods. SpinQuant (Liu et al., 2024c) optimizes full $n\times n$ rotation matrices on the Stiefel manifold, requiring $O(n^{2})$ parameters (16.7M for 4096-dim layers) and $O(n^{3})$ operations. ROSAQ (Yoon et al., 2025) and KurTail (Akhondzadeh et al., 2025) learn saliency-aware and kurtosis-guided rotations respectively, but lack theoretical guarantees and still require $O(n^{2})$ space. DuQuant (Lin et al., 2024a) employs a greedy construction method that uses identified outlier dimensions as prior knowledge to build rotation matrices, followed by zigzag permutation to balance outlier distribution across blocks.
Our Position. ButterflyQuant bridges fixed and learned approaches through structured parameterization. Unlike QuIP#’s fixed rotations or SpinQuant’s expensive Stiefel optimization, our Givens parameterization guarantees orthogonality by construction while enabling efficient data-driven learning, combining theoretical guarantees with adaptability. Crucially, the butterfly computational graph is structurally identical to the fast Hadamard transform, enabling adaptation of existing optimized inference kernels with minimal modifications.
@@ 2.3 Structured Transforms and Orthogonal Parameterization
Despite extensive use in deep learning, structured transforms remain unexplored for quantization, which requires: (1) strict orthogonality for computational invariance $\mathbf{y}=\mathbf{Wx}=(\mathbf{WQ}^{T})(\mathbf{Qx})$, (2) adaptability to layer-specific outliers (Sun et al., 2024a), and (3) efficient inference. While FlatQuant (Sun et al., 2024b) uses Kronecker decomposition, it sacrifices orthogonality. ButterflyQuant uniquely combines continuous parameterization ($\theta\in\mathbb{R}$ enabling gradients vs. Hadamard’s discrete $\{+1,-1\}$) and $O(n\log n)$ complexity, balancing expressiveness with efficiency. ButterflyQuant is the pioneer in leveraging these properties for LLM quantization.
@@ 3.1 Preliminaries: Rotation-Based Quantization and Incoherence
Given a weight matrix $\mathbf{W}\in^{m\times n}$ and activation $\mathbf{x}\in^{n}$, standard quantization directly quantizes:
@@ 3.1.1 Theoretical Foundation: The Incoherence Principle
The key insight from QuIP (Chee et al., 2023) is that quantization error is minimized when the rotation basis is maximally incoherent with the standard basis. For an orthogonal matrix $\mathbf{Q}\in\mathbb{R}^{n\times n}$, the mutual coherence is defined as:
@@ 3.1.2 Fixed Hadamard Transforms
QuaRot (Ashkboos et al., 2024) applies Hadamard transforms as a computationally efficient approximation to random rotations. The Hadamard matrix $\mathbf{H}_{n}$ of dimension $n\times n$ is recursively defined as:
@@ 3.1.3 From Fixed to Learnable Continuous Rotations
Although Hadamard transforms achieve optimal worst-case incoherence, they suffer from two critical limitations: (1) Their discrete $\{+1,-1\}$ entries prohibit gradient-based optimization, and (2) neural networks exhibit layer-specific structured patterns (Sun et al., 2024a). As shown in Figure 1(a), attention layers, early MLPs, and late MLPs each have distinct outlier distributions that require tailored rotations. Fixed transforms cannot adapt to these heterogeneous patterns, treating all layers identically despite their vastly different quantization challenges.
@@ 3.2.1 Structure and Parameterization
This parameterization ensures orthogonality by construction while enabling gradient-based optimization. See Appendix 7 for detailed matrix structures and visualizations.
@@ 3.2.2 Relationship to Hadamard Transforms
Theorem 3.1.
The Hadamard matrix $\mathbf{H}_{n}$ for $n=2^{k}$ can be exactly represented as a butterfly transform with specific parameter choices (Dao et al., 2019).
@@
$\mathbf{H}_{n}$ $=\begin{bmatrix}\mathbf{H}_{n/2}&\mathbf{H}_{n/2}\\ \mathbf{H}_{n/2}&-\mathbf{H}_{n/2}\end{bmatrix}$ (10)
$=\begin{bmatrix}\mathbf{I}_{n/2}&\mathbf{I}_{n/2}\\ \mathbf{I}_{n/2}&-\mathbf{I}_{n/2}\end{bmatrix}\begin{bmatrix}\mathbf{H}_{n/2}&\mathbf{0}\\ \mathbf{0}&\mathbf{H}_{n/2}\end{bmatrix}$
@@
which corresponds to a $2\times 2$ orthogonal matrix with angle parameter $\theta=\pi/4$. Specifically, $\mathbf{H}_{2}=\mathbf{G}(\pi/4)\cdot\text{diag}(1,-1)$, where the diagonal sign matrix is itself orthogonal and can be absorbed into the butterfly factorization. More generally, each $2\times 2$ block in the Hadamard factorization is a product of a Givens rotation and a diagonal sign matrix, both of which are orthogonal and continuously parameterizable. The complete recursive factorization yields $\log_{2}n$ sparse matrices, each implementable as a butterfly layer with specific angle and sign choices, establishing that Hadamard transforms are a special case of butterfly parameterizations. ∎
@@ 3.2.3 Theoretical Coherence Analysis
For orthogonal transforms, coherence $\mu(\mathbf{Q})=\max_{i\neq j}|Q_{ij}|$ measures how evenly information is distributed. From compressed sensing (Candès et al., 2006; Donoho, 2006), lower coherence reduces the sampling requirements for successful recovery:
@@
Figure 2 compares coherence across LLaMA-2-7B layers ($n=4096$). Hadamard transforms achieve the Welch bound $\mu(\mathbf{H}_{n})=1/\sqrt{n}=1.56\times 10^{-2}$ uniformly. Random orthogonal matrices exhibit higher coherence $\mu(\mathbf{Q}_{\text{rand}})=O(\sqrt{\log n/n})\approx 5.4\times 10^{-2}$ (Vershynin, 2018). Learned butterfly transforms demonstrate adaptive coherence varying from $1.8$ to $3.2\times 10^{-2}$ across layers, matching the heterogeneous outlier patterns: early attention layers maintain near-Welch-bound coherence for uniform decorrelation, while deeper MLP layers relax this constraint for their specific activation patterns.